OpenStudy (anonymous):

sqrt[3]{80}+9sqrt{125}

OpenStudy (anonymous):

\[\sqrt{3}*80 + 9\sqrt{125}\]

OpenStudy (anonymous):

\[\sqrt[3]{80}+9\sqrt{125}\]

OpenStudy (anonymous):

??

OpenStudy (alexwee123):

 or is 80 out of the radical?